OpenStudy (anonymous):

Evaluate the following indefinite integral \[\int \frac{1+x}{1+x^{2}}dx\] Transforming it to a familiar form, we get; \[\int \left ((1+x)\cdot \left ( \frac{1}{1+x^{2}} \right ) \right )dx\] From here I cannot find the proper substitute to solve the problem

OpenStudy (anonymous):

perform integration by parts. might help you.

OpenStudy (anonymous):

I could do that, but the work is on substitution :x

OpenStudy (nikvist):

\[\int udv=uv-\int vdu\quad;\quad u=1+x\quad,\quad dv=\frac{dx}{1+x^2}\]